OpenStudy (anonymous):

(Optimization) Find the volume of the largest right circular cylinder that can be inscribed in a sphere of radius r.

OpenStudy (anonymous):

Um... @hartnn , can u please help me. x)

hartnn (hartnn):

could u form an equation ?

OpenStudy (anonymous):

V(sphere)=(4/3)pi(r)^3 and V(cylender)=pi(r)^2h is this what u mean?

OpenStudy (anonymous):